# Applying Text Embedding Models for Efficient Analysis in Labeled Property Graphs

### v1

@@ 1 Introduction
Labeled property graphs (LPGs) represent a flexible data model in which nodes and edges are not only connected through relationships but also enriched with key-value properties, many of which are textual. This structure is well-suited to domains such as social networks, knowledge graphs, and institutional data, where descriptive attributes carry important semantic information. Despite this, most analytical approaches to LPGs focus primarily on structural information - such as connectivity patterns or edge types - while treating textual fields as flat labels or ignoring them altogether.
@@
In this work, we investigate how text embedding models can be used to enhance property graph analysis. Specifically, we apply them to two fundamental tasks: node classification and link prediction. In both cases, textual node and edge attributes are transformed into semantic vector representations using a state-of-the-art embedding model. These embeddings are then used as inputs to lightweight classifiers, enabling the graph to be analyzed through its language rather than just its structure.
We evaluate this approach on Neo4j [1] real-world example graph datasets from diverse domains, demonstrating that pretrained text embeddings alone can yield strong predictive performance. The results suggest that integrating textual semantics into LPG workflows is a practical and scalable strategy for improving the accuracy and interpretability of graph-based analysis.
@@ 2 Related work
Recent work at the intersection of graphs and language includes combining structural and textual features in knowledge graphs [7, 8], but these methods often assume fixed schemas or require joint training. Property graphs, with their flexible labeled attributes, remain underexplored in this context.
Our approach differs by integrating pretrained text embedding models into property graph analysis without modifying the graph structure. This enables semantically informed classification and relation prediction using existing text fields as-is.
@@ 3 Datasets
We use two publicly available labeled property graph datasets provided by Neo4j, a widely adopted graph database platform. These datasets represent diverse domains and graph structures:
  • Twitter Trolls [13]: A social network graph capturing interactions between accounts linked to coordinated disinformation campaigns.
  • WWC 2019 [12]: A sports-focused dataset modeling the 2019 FIFA Women’s World Cup, including players, teams, matches, and events.
These datasets offer rich textual properties on nodes and relationships, making them well-suited for evaluating the integration of text embedding models into graph analysis tasks such as node classification and relation prediction.
@@ 4 Solution
In the first task, we aim to classify nodes based on their textual attributes. For each node, we extract available properties and convert them into a coherent textual representation. This input is passed through the Qwen3 embedding model, producing a 1024-dimensional vector that encodes the semantic content of the node. These embeddings are then used as input features for a standard classifier, such as logistic regression or a random forest, trained to predict labels of the nodes. The approach supports generalization to unlabeled nodes and provides a lightweight, language-model-powered alternative to structural graph embedding methods.
The second task involves relation prediction, where the goal is to infer missing edges between node pairs. For this, we construct a more detailed textual prompt describing a source node, its neighbors and associated edge types, and optionally the target node. This full context is serialized into a single string and embedded using Qwen3. The resulting vector captures both local relational structure and semantic content. A classifier is trained to predict the likely relation type between the source and target nodes based on this embedded context. During training, existing edges are removed to serve as prediction targets, and randomly sampled node pairs without a relation are included as negative examples to improve discriminative power.
@@ 5 Results
The proposed approach was applied to two labeled property graph datasets from Neo4j - Twitter Trolls and WWC 2019 - each containing varied node types and rich textual attributes. In the node classification task, embeddings generated by the Qwen3-Embedding-0.6B model enabled accurate prediction of categories such as player affiliations or entity roles, using simple classifiers trained directly on the semantic vectors. For relation prediction, the method successfully recovered removed edges, based solely on node descriptions and local context. The results were consistently strong across datasets, with high classification accuracy and clearly meaningful relation predictions. The model handled both sparse and dense text fields effectively and showed robustness to inconsistencies or missing attributes. Overall, the combination of text embeddings and labeled property graph structure proved to be an effective and efficient foundation for semantic graph analysis.
Table 1 shows classifier performance for predicting REPRESENTS relations between players and countries in the WWC 2019 dataset.
@@
k-Nearest Neighbors 0.976 0.977 0.976 0.975
Table 1: Classifier performance for predicting REPRESENTS relations between players and countries in the WWC 2019 dataset.
Table 2 shows classifier performance node labels in the WWC 2019 dataset.
@@
Random Forest 0.992 0.992 0.992 0.990
Logistic Regression 0.993 0.993 0.993 0.991
SGDClassifier 0.995 0.995 0.995 0.994
Support Vector Machine 1.000 1.000 1.000 1.000
k-Nearest Neighbors 0.997 0.997 0.997 0.997
Table 2: Classifier performance for node labels in the WWC 2019 dataset.
Table 3 shows classifier performance for predicting node labels in the Twitter Trolls dataset.
@@
Random Forest 0.984 0.984 0.984 0.983
Logistic Regression 0.999 0.999 0.999 0.999
SGDClassifier 0.994 0.993 0.994 0.994
Support Vector Machine 1.000 1.000 1.000 1.000
k-Nearest Neighbors 0.989 0.989 0.989 0.989
Table 3: Classifier performance for predicting node labels in the Twitter Trolls dataset.
@@ 6 Next Steps
Building on these initial results, several directions offer potential for further development. One priority is to explore task - specific fine-tuning of the embedding model, particularly in domains with specialized language or metadata. Incorporating structural information more explicitly - through hybrid models that combine text embeddings with graph neural networks - may also improve performance in tasks where topology plays a stronger role. Another avenue is to extend the current approach to enable embedding-based semantic search and similarity-driven querying directly within the graph environment. Additionally, scaling the method to larger graphs will require optimizing embedding workflows, including caching strategies and batch processing. Finally, applying this approach to more complex downstream tasks such as community detection, role inference, or temporal relation analysis could demonstrate its broader applicability across graph-based systems.
@@ 7 Conclusions
This method is model-agnostic, lightweight, and compatible with existing property graph platforms. Our evaluation on Neo4j datasets demonstrates its broad applicability across domains.

### v2

@@ 1 Introduction
Labeled property graphs (LPGs) represent a flexible data model in which nodes and edges are not only connected through relationships but also enriched with key-value properties, most of which are textual [1]. This structure is well-suited to domains such as knowledge graphs, social networks, and institutional data, where descriptive attributes carry important semantic information. Despite this, most analytical approaches to LPGs focus primarily on structural information - such as connectivity patterns or edge types - while treating textual fields as flat labels or ignoring them altogether.
@@
In this work, we investigate how text embedding models can be used to enhance property graph analysis. Specifically, we apply them to two fundamental tasks: node classification and link prediction. In both cases, textual attributes are transformed into semantic vector representations using a state-of-the-art embedding model. These embeddings are then used as inputs to lightweight classifiers, enabling the graph to be analyzed through its language rather than just its structure.
The results suggest that integrating textual semantics into LPG workflows is a practical and scalable strategy for improving the accuracy and interpretability of graph-based analysis.
@@ 2 Related work
Recent work at the intersection of graphs and language includes combining structural and textual features in graphs [7, 8], but these methods often assume fixed schemas or require joint training. Property graphs, with their flexible attributes, remain underexplored in this context.
Our approach differs by integrating pretrained text embedding models into labeled property graph analysis without modifying the graph structure. This enables semantically informed classification and relation prediction using existing text fields as-is.
@@ 3 Datasets
We use publicly available labeled property graph datasets provided by Neo4j [9], a widely adopted graph database platform. Used datasets represent diverse domains and graph structures:
  • Twitter Trolls (nodes 281136, edges 493160) [10]: A social network graph capturing interactions between accounts linked to coordinated disinformation campaigns.
  • Legis (nodes 11825, edges 523004) [11]: A knowledge graph representation of the U.S. Congress comprising legislators, bills, committees, votes, and related entities.
  • WWC 2019 (nodes 2486, edges 14799) [12]: A sports-focused graph dataset modeling the 2019 FIFA Women’s World Cup, including players, teams, matches, and events.
  • Stack Overflow (nodes 6193, edges 11540) [13]: A graph modeling Stack Overflow questions, answers, tags, comments, and the relationships between them.
These datasets offer rich textual properties on nodes and relationships, making them well-suited for evaluating the integration of text embedding models into graph analysis tasks.
@@ 4 Solution
We formulate two prediction tasks based on language-model embeddings of graph nodes. In both cases, the available textual and relational information for node is serialized into a coherent string and processed by the embedding model to produce a 1024-dimensional vector representation. Instances are randomly split into training and test sets, with 90% used for training and 10% reserved for evaluation. Standard classifiers - Random Forest, Logistic Regression, SGDClassifier, and Support Vector Machine (SVM) - are trained on the embeddings and evaluated on the held-out test set.
In the node label prediction task, the input string contains the textual properties of the node, combined into a single coherent description. The classifiers learn to map these embeddings to the corresponding node labels and can subsequently be applied to unlabeled nodes.
In the relation prediction task, the goal is to infer missing factual information (Figure 1). For each labeled instance, a specific relation of the source node is withheld (for example, the relation between a player and a team). The remaining relations of the source node, together with its connected neighbor nodes, are included in the textual description. The trained classifiers predict the correct target node, thereby recovering the withheld relation.
Figure 1: Relation Prediction Setting: In this setting, a specific relation of the source node (e.g., between a Player and a Team in the WWC 2019 dataset) is withheld prior to embedding generation. The embedding model encodes the source node based on its remaining relations and neighboring nodes. The resulting embedding is then used in a downstream classification task to predict the correct target node, thereby recovering the withheld relation.
@@ 5 Results
The proposed approach was applied to property graphs containing varied node types and rich textual attributes. In the node classification task, generated embeddings enabled accurate prediction of labels. Results are presented in Table 1. For relation prediction, the method successfully recovered information carried by removed edges, based solely on node descriptions and local context. Representative results are gathered in Table 2.
The results were consistently strong across datasets, with high classification accuracy and clearly meaningful relation predictions. The model handled both sparse and dense text fields effectively and showed robustness to inconsistencies or missing attributes. Overall, the combination of text embeddings and labeled property graph structure proved to be an effective and efficient foundation for semantic graph analysis.
Table 1: Classifier performance for predicting node labels across datasets.
Classifier Accuracy Precision Recall F1 Score
Random Forest 0.992 0.992 0.992 0.990
Logistic Regression 0.993 0.993 0.993 0.991
SGDClassifier 0.995 0.995 0.995 0.994
Support Vector Machine 1.000 1.000 1.000 1.000
(a) WWC 2019
Classifier Accuracy Precision Recall F1 Score
Random Forest 0.984 0.984 0.984 0.983
Logistic Regression 0.999 0.999 0.999 0.999
SGDClassifier 0.994 0.993 0.994 0.994
Support Vector Machine 1.000 1.000 1.000 1.000
(b) Twitter Trolls
Classifier Accuracy Precision Recall F1 Score
Random Forest 0.822 0.823 0.822 0.821
Logistic Regression 0.926 0.926 0.926 0.926
SGDClassifier 0.928 0.930 0.928 0.929
Support Vector Machine 0.939 0.939 0.939 0.939
(c) Stack Overflow
Table 2: Classifier performance for link prediction across datasets.
@@
(a) WCC 2019 - REPRESENTS relation
@@
Random Forest 0.796 0.807 0.796 0.786
Logistic Regression 0.858 0.875 0.858 0.851
SGDClassifier 0.919 0.914 0.919 0.916
Support Vector Machine 0.845 0.862 0.845 0.837
(b) Legis - IS_MEMBER_OF relation
@@
Random Forest 0.987 0.987 0.987 0.987
Logistic Regression 0.993 0.993 0.993 0.993
SGDClassifier 0.987 0.987 0.987 0.987
Support Vector Machine 0.987 0.987 0.987 0.987
(c) Legis - ELECTED_TO relation
@@ 6 Next Steps
Future work could focus on enhancing the integration of textual and structural information, for example by combining language-model embeddings with topology-aware graph encoders. Another promising direction is task-specific adaptation of the embedding space to capture domain-specific semantics and relation patterns. Scaling to large and dynamic graphs may require efficient incremental embedding updates and distributed processing pipelines. Finally, systematic evaluation across heterogeneous datasets would help to assess generalization and identify domain-dependent optimization strategies.
@@ 7 Conclusions
This method is model-agnostic, lightweight, and compatible with existing property graph platforms. Our evaluation demonstrates its broad applicability across domains.